OpenStudy (anonymous):

the sum of three consecutive numbers is less than or equal to 216. find the greatest possible values of these integers

jimthompson5910 (jim_thompson5910):

let x be the first number so x+1 is the next number x+2 is the next one after that

jimthompson5910 (jim_thompson5910):

the sum of three consecutive numbers ---> (x)+(x+1)+(x+2) ---> 3x + 3

jimthompson5910 (jim_thompson5910):

so... the sum of three consecutive numbers is less than or equal to 216 translates to 3x + 3 <= 216

jimthompson5910 (jim_thompson5910):

solve for x and tell me what you get

OpenStudy (anonymous):

i got 71 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

x <= 71 is correct so 71 is the largest value x can be

jimthompson5910 (jim_thompson5910):

which means x+2 = 71+2 = 73 is the largest value possible

OpenStudy (anonymous):

so would i write 71 down or 73 as my answer @jim_thompson5910

jimthompson5910 (jim_thompson5910):

73 since that's the larger of the two

jimthompson5910 (jim_thompson5910):

actually I think they want all 3, so 71, 72, 73

jimthompson5910 (jim_thompson5910):

they want the "greatest possible values of these integers" so you're listing them all (the largest they can be)

OpenStudy (anonymous):

ok thanks ... can you help with this one the sum of three consecutive odd numbers is more than 207. find the minimum values of these integers @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the sum of three consecutive odd numbers is more than 207 translates to (2x+1)+(2x+3)+(2x+5) > 207 solve for x

OpenStudy (anonymous):

i got 33 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

x > 33, yep

jimthompson5910 (jim_thompson5910):

now plug in x = 34 (the next integer after 33) 2x+1 = 2*34+1 = 69 2x+3 = 2*34+3 = 71 2x+5 = 2*34+5 = 73

jimthompson5910 (jim_thompson5910):

so the smallest 3 integers are: 69, 71, 73